OpenStudy (anonymous):

Simplify (6/x)/(6/x+1)

OpenStudy (anonymous):

x/(x+1)

OpenStudy (anonymous):

\[\frac{\frac{6}{x}}{\frac{6}{x+1}}=\frac{6}{x}*\frac{x+1}{6}=\frac{6(x+1)}{6x}=\frac{(x+1)}{x}=1+\frac{1}{x}\]

OpenStudy (anonymous):

sorry first post was incorrect

OpenStudy (anonymous):

says its wrong :(